### v1

# Anatomy of Elite and Mass Polarization in Social Networks

Abstract
Existing methods for quantifying polarization in social networks typically report a single value describing the amount of polarization in a social system. While this approach can be used to confirm the observation that many societies have witnessed an increase in political polarization in recent years, it misses the complexities that could be used to understand the reasons behind this phenomenon. Notably, opposing groups can have unequal impact on polarization, and the elites are often understood to be more divided than the masses, making it critical to differentiate their roles in polarized systems. We propose a method to characterize these distinct hierarchies in polarized networks, enabling separate polarization measurements for these groups within a single social system. Applied to polarized topics in the Finnish Twittersphere surrounding the 2019 and 2023 parliamentary elections, our analysis reveals valuable insights: 1) The impact of opposing groups on observed polarization is rarely balanced, and 2) while the elite strongly contributes to structural polarization and consistently display greater alignment across various topics, the masses have also recently experienced a surge in issue alignment, a special form of polarization. Our findings suggest that the masses may not be as immune to an increasingly polarized environment as previously thought.
1 Introduction
Distinct aspects of political polarization are reportedly higher today in many societies [ILL+19, Sil18, KYA21, FZQ+23, BG20, KC24]. It is a serious and consequential phenomenon that has garnered significant attention in recent years [FGT+22, XGM+22, FGF+23, Kaw22]. It refers to the increasing divergence of political attitudes, beliefs, and values between individuals and groups [DEB96, FA08], often leading to heightened hostility [FH23, BC18], ideological rigidity [WCY+22], and a breakdown in constructive dialogue [FH23]. As polarization intensifies, it poses significant challenges to democratic societies [MRS18], hindering effective governance [Jon01] and compromising social cohesion [Mas15]. Hence the implications of political polarization are far-reaching, affecting policy-making [WHD+21], public discourse [RM16], and even personal relationships [WCP21, GS23]. Understanding the causes, dynamics, and consequences of political polarization is crucial in order to address its negative impacts.
Current methods for quantifying polarization in social networks assign a single score to an entire social system (e.g., [GMGM18, GMJCK13, MBLB15, LAD24, GAS+15]), simplifying the complexity of the underlying phenomenon. While these methods are often intuitive and straightforward to implement, they produce high polarization values for random networks [SCK22] and fail to capture many crucial nuances related to the multifaceted nature of polarization. They do not distinguish between the elite and the mass polarization, nor do they reveal whether opposing groups contribute equally to the observed division in a given network. These methods are also typically designed to measure the decreasing communication between groups and/or increasing cohesion within them, thus missing other aspects of polarization, such as political alignment [Mas15, Tör22, CSG+21].
We argue that knowing whether both conceptually and structurally dominant individuals are more polarized than the less dominant individuals is important for understanding such polarized systems. Moreover, describing overall polarization with a single score can create the false perception that all groups are equally influential in the observed polarization online. Despite a growing body of work reporting increasing polarization worldwide [AR21, CH23, KYA21, FGT+22, KC24, BG20], we lack standardized methods to break down the measured division into more interpretable components that provide deeper insights into the phenomenon. In this study, we aim to extend the informativity of common polarization measures. We begin by identifying the assortative clusters [ZP20] representing the polarized groups. Then, we further partition each identified group into a hierarchy using the well-known core-periphery structure [BE00, RPFM17, GYW21] to separate the elite from the mass. Finally, we study two distinct aspects of polarization: structural polarization and issue alignment, at both the group and hierarchical levels. For the former, this involves decomposing an existing polarization measure, while for the latter, we measure the degree of alignment separately for the identified hierarchical groups.
Some polarization studies focus solely on the political elite (e.g. [Nea20]), such as elected officials, media figures or influential individuals, while some are more limited to assess the division among the general mass (e.g. [FA08, Lel16]). Researchers have also explored the interplay between these two groups (elite and mass) in polarized settings (e.g. [BC18, BCRR23, Sky21, RM16, FA08, LC02]), resulting in valuable insights into their differing levels of political responsiveness [BCN+19] and ideological consistency [LC02]. Generally, the political elite are perceived as exhibiting greater ideological coherence and predictability in their positions across various domains compared to the mass. Some theories (e.g. [LC02, KK17]) suggest that the public are actually immune to elite polarization, whereas political elites in the opposing groups, are growing more ideologically distant from each other while also becoming more internally homogeneous [RM16]. However, recent body of work has found people’s views converging and becoming more interwined, possibly indicating that certain ideological realignment is taking place among the mass too [KM21]. To our knowledge, prior research on distinguishing the contribution of the elite and the mass to distinct aspects of polarization in online systems is rather limited. Our approach will help us evaluate whether polarization among the mass has remained at the same level, independent of the polarization trends observed among the elite.
Additionally, there exists work underscoring the importance of reconsidering the idea of treating both sides of a political discussion as equal [JBD22, GR14, SRZ19]. The language [KC24], rhetoric, and distribution of power can vary significantly between different ideological groups. Equivalently, polarization can be also asymmetrical [JBD22], which indicates that a group can become more extreme while the other group staying at the same level. The majority of current methodologies struggle to discern which group or groups contribute more significantly to overall structural polarization or issue alignment. Implementing a consistent partition strategy with an intuitive characterization of identified structures allows for a more precise description of disparities between these polarized groups and enables comparisons across different topics represented by networks.
Generally, assuming that each individual or group contributes uniformly to polarization can be suboptimal for several reasons. As the literature has shown, we should expect that the nature of an individual’s divisiveness depends on their role in society. We also know that the mass can take social and political cues from political elites [VBS21, DL19, GEN+20, Sky21], and are often influenced by them [All23, MS21, KT18, GEN+20, Ber07, MS18]. When elites indicate a divergence in ideology or display hostility towards out-groups, their followers interpret these signals and may adjust their behavior accordingly [BC18, FH23, Ber07, MS18]. This phenomenon, commonly referred to as elite signaling or elite influence, has been linked to political polarization [BCRR23, Sky21]. While the impactful role of elite in political and social systems is evident, one should not underestimate the impact of the mass either. Previous research suggests that politically-engaged individuals within the public may be the ones setting the political agendas for discussions on online platforms [BCN+19]. Not only might the people be leading the conversations, they are often needed for reaching wider audience in the social networks [BWB+15]. While the connection between these unique characteristics of the elite and mass groups and polarization is complex and, for the most part, unknown, it is crucial to examine how they differ in their contribution to the political polarization observed online.
In recent decades, three distinct aspects of polarization have received significant attention in the community: 1) the bimodality of opinion distributions, 2) the extent of issue and partisan alignment, and 3) the affective dimension, which refers to the development of strong negative attitudes and emotions towards members of opposing political or ideological groups. While these three aspects of polarization can potentially strengthen each other, it is important to recognize that they are different concepts, both theoretically and in terms of how they are measured empirically [HDC23]. In this paper, we focus on the first two aspects.
Researchers have traditionally examined the bimodality and alignment aspects of polarization using survey-based data, estimating ideological shifts and correlations between distinct issue positions from participants’ responses (e.g. [KM21]). However, the research community has increasingly turned its attention to social media platforms, as, in today’s world, these platforms serve as crucial venues and spaces for daily discussions on various political topics and debates. Additionally, these online platforms have become arenas for political competition, with millions of interactions and users, leading to valuable insights into the ideological distances among the users, their limited communication, and elite influence on public opinion worldwide (e.g. [Bri18, CRF+11, CMK+20, GEN+20]). Social media platforms enable researchers to monitor conversations and interactions in almost real-time, allowing for timely analysis of emerging trends and shifts in public opinion. Besides the inexpensive access to large datasets, it is also reasonable to expect that there are less response bias in users’ online behavior compared to traditional survey experiments. These advantages may explain the growing interest in analyzing polarization through the structural properties of networks.
Within social networks the bimodality of opinion distribution can be estimated by measuring the limited communication or reachability between the opposing groups in a given network [MBLB15], although some recent work have advocated for handling these separately [HDC23]. The intuition is that sharper and further peaks in an opinion distribution related to a topic X would correspond to a more isolated communities in a network where the topic X is being discussed. This type of isolation is often referred to as structural [SCK22] or interactional [FZQ+23] polarization. Most of the methods designed to capture this assesses the separation of interaction patterns among two distinct groups [GMGM18] — although some work have been recently done for handling multiple groups [NI22, MGLB23]. Typically, quantifying the structural polarization of a political communication network involves a three-step process [GMGM18, SCK22]. Firstly, the network is constructed based on the collected data. Following this, community detection techniques are employed to identify functional groups, often referred to as polarized groups, within the network, which frequently represent opposing sides of a political discourse. Finally, the quality of the division is evaluated using some structural polarization score, such as Random Walk Controversy [GMGM18] or Adaptive EI-index [CSG+21]. Again, the assumption these methods often have is that this division between the two clusters corresponds to either the lack of interactions between the clusters or the degree to which individuals are trapped within their own communities. Moreover, many of these methods require identifying the key figures in a network in order to quantify the extent of separation [GMGM18, SCK22, MBLB15], in which typically researchers rely on the number of endorsements a user has as a proxy for their importance or “eliteness”. While some polarization studies utilize known node labels [FGT+22, XGM+22] to identify e.g. politicians, this information is not always available and might miss other crucial users without clear affiliations, such as nonpolitical opinion leaders, which are shown to be more popular than the political figures on social media [MJL22]. Although our method is not specifically designed to address this potential limitation, there are situations where inferring the key figures in a network solely from certain structural properties may be preferable.
While these methods have produced valuable insights into societal division trends, their approach to measuring polarization may be somewhat restrictive. In addition to the inflexibility in choosing the number of groups, they are primarily designed to quantify the extent of interactional patterns within a single issue, and hence, not being capable of capturing the potential correlation between individuals’ stances on different topics. Some political and social polarization literature identifies issue alignment as more deleterious aspect of polarization [Mas15, CSG+21, Tör22]. Issue alignment refers to the degree of agreement or similarity between individuals or groups on a set of issues. In issue alignment, individuals or groups may align themselves based on shared values or ideologies leading to a collective stance on range of issues, potentially deepening divisions in the society [Mas15, Tör22].
Our contribution comprises both methodological and exploratory aspects. We study five topics that have become drastically more polarized online in Finland over the course of four years, spanning the period from 2019 to 2023, which contains major events, such as a global pandemic and a war in Europe. In this paper, we apply this framework to analyze polarization trends and highlight their differences between two snapshots of the Finnish Twittersphere. Specifically, we will examine the periods surrounding two parliamentary elections, in 2019 and 2023, and demonstrate how polarization dynamics vary across networks. We find that the polarized groups, as well as the elite and mass within them, shape polarization trends in very different ways.
To reiterate, our objective is to propose a method capable of measuring two distinct types of polarization—structural polarization and issue alignment—within both elite and public spheres, utilizing network representations of social or political systems. Additionally, our method unravels the imbalances in the impact that distinct polarized groups have on the observed polarization—something that a single polarization value would not reveal.
This paper is organized as follows: In Section 2, we outline the steps for inferring hierarchical structures in a polarized network and introduce our methods, which utilize these hierarchies to describe polarization. The main findings are discussed in Section 3, where we study the trends and patterns in hierarchical polarization on Twitter during the Finnish parliamentary elections in 2019 and 2023. Finally, in Section 4, we conclude with a discussion, list of identified limitations and ideas for future work.
2 Hierarchies in polarized networks
Figure 1: Observed network is depicted by the gray block, and it is divided into two assortative groups (pink and turquoise blocks) only when there is sufficient evidence of such groupings in the network. Once the polarized groups are identified, each block is further partitioned into a core and periphery independently, given significance of the hierarchy. Following the hierarchy decomposition, there are four potential scenarios: both polarized groups exhibit a hierarchical structure (1), one of the groups demonstrates a hierarchical structure (2-3), or neither of the groups displays indications of a hierarchical structure (4).
2.1 Finding the groups and hierarchies
Polarization is often assumed to significantly impact social network structures, giving rise to two internally tightly-knit groups with sparse connections between them [GMGM18, HDC23]. Therefore, the level of polarization in a system can be quantified by examining the patterns of interactions within that configuration [FGT+22, GMGM18]. To conduct our experiments, our initial step is to identify the polarized groups. Following this, we need to distinguish between the elite and the mass members within each polarized group. This means that each individual in a network belongs to a specific polarized group (for-against) and a hierarchical group (elite-mass). While from now we will focus on two polarized groups for this study, this method can be readily adapted to scenarios involving more than two opposing groups.
Decomposing a network involves breaking it into modular and structurally distinct components. The conventional approach typically entails identifying the polarized groups in a network using clustering algorithms [GMGM18]. However, prior research has highlighted that many clustering algorithms can generate convincing group structures even in entirely random networks [LRR10, GSPA04, ZM14]. This tendency results in non-controversial networks having inflated polarization values, leading to a high number of false positives [SCK22]. To alleviate this issue, we employ the stochastic block model (SBM) to identify both polarized groups and hierarchical groups in a network, exploiting the clustering method’s ability to compute the description length for the given model and partition. Minimum description length is a principle for model selection that favors models with the shortest code length needed to describe the data, consistent with Occam’s razor, which suggests simpler explanations are generally preferable [Grü07]. Thus, an observed community structure that could be explained due to randomness, would be flagged with SBM, as the description length for an SBM with no community structure would be lower than that of a model attempting to overfit by dividing the network into two communities [Yan16, Pei19, Pei14b, ZP20]. Note, however, that this does not completely eliminate the need of normalizing the polarization scores for instance in classification tasks [SCK22].
There are two assumptions that are translated into our choices of SBM models. According to our first assumption a polarized network consists of two assortative groups, and thus, we apply the planted partition model variation of the SBM [ZP20, Pei14b]. We sweep the number of blocks (groups) between 1 to 2, and select the configuration leading to the lowest description length.
Our second assumption is the existence of hierarchies in each polarized group. Social groups tend to naturally self-organize into hierarchies, where members exhibit varying levels of influence, expertise, or dominance [KXO15]. The emergence of these hierarchies is considered both inevitable and often advantageous for social cohesion [KXO15, MG08]. We utilize the well-established concept of core-periphery structures (see e.g. [RPFM17]) in approximating the hierarchical nature of these polarized groups. The core-periphery structure is a common pattern observed in various systems, including social networks [BE00, KM17, YZS+18], economics [CLWU13, Wan16], and many other fields [CLWU13, KM17]. It consists of a densely interconnected core, typically representing nodes with higher connectivity, influence, or resource access, surrounded by a sparsely connected periphery representing nodes with lesser influence. In our context, the core is equivalent to the elite, and the periphery is equivalent to the mass (or public).
We anticipate that the elite are typically positioned within structurally dominant areas of networks, forming cohesive structures [MJRW20]. On Twitter, the elite often include influential figures such as politicians, journalists, activists, and experts who are active on social media. In contrast, we assume the surrounding periphery comprising the mass, consisting primarily of users who endorse or amplify the core’s voice [BWB+15]. These assumptions are later assesed for our data in Section 3.
For identifying the hierarchical structures that are embedded in the already identified polarized groups, we follow the method proposed by Gallagher et al. to infer the core-periphery structure [GYW21] within the SBM framework. We treat each group independently, which enables us to compare their observed hierarchies to the null model separately. We use the common Erdős–Rényi model as the baseline, as done for instance in [KM18], because it assumes no inherent hierarchy, even one that might be explained by the network’s degree sequence. This approach has its pros and cons that we address separately in the Appendix A.7. A visual depiction of the overall network decomposition pipeline is shown in Figure 1. Note that alternative methods for identifying core-periphery structures exist, such as the extended version of the original method [BE00] for finding multiple core-periphery pairs [KM17] or the method based on detecting core-periphery structures by surprise [dJCS19].
(a) Full network
(b) Cores
(c) Peripheries
Figure 2: The complete retweet network representing Climate discussion (2023) in (a), subgraph containing the identified cores in (b), and subgraph containing peripheries in (c). The complete observed graph is drawn with ForceAtlas2 -layout algorithm [JVHB14]. For the others graphs, same node and edge positions are applied with specific hierarchy groups filtered out.
2.2 Defining structural polarization and issue alignment for hierarchies
Prior work has shown that the most common quantitative methods for measuring structural polarization in social networks perform similarly after they are denoised [SCK22]. According to the same work, the best performing method in detecting polarized networks was Adaptive EI-index (AEI), which is based on the more popular EI-index (e.g. [EDVBP22, Yan16, DVB18, Bor23]) developed by Krackhardt and Stern [KS88]. This method produces a score based on the distribution of connections among nodes within and between groups, where a greater density of links within groups relative to those between groups indicates a more polarized system.
Let $G=(V,M)$ represent a network, where $V$ is the set of nodes and $M$ is the set of links. We define a partition of $G$ into two groups, denoted as $A$ and $B$, such that: $V=V_{A}\cup V_{B}$ and $V_{A}\cap V_{B}=\emptyset$, where $V_{A}$ and $V_{B}$ represent the sets of nodes belonging to groups $A$ and $B$ respectively. The AEI is defined for such a system as
$P_{AEI}=\frac{i_{A}+i_{B}-2\times e_{AB}}{i_{A}+i_{B}+2\times e_{AB}}\,,$ (1)
where $i_{X}$ denotes the internal links density (observed links divided by the possible number of links) within group $X$ and $e_{AB}$ denotes the external links density between groups $A$ and $B$.
In the AEI-index, all links are treated homogeneously within each group, which prevents it from distinguishing the activity driven by the elite versus the mass. However, one should not view these links equal as the elite are often seen as highly influential individuals on social platforms due to their large following, prominent status in society and access to resources that can amplify their messages and shape the public opinion. An elite individual can also connect to other elite members enabling them to reach even a wider audience and have a greater impact on conversations and trends within their communities, making them “more organized”. The amplification and social validation come mostly from the followers, the surrounding mass, who both take cues from the elite and can be mobilized by them to take specific actions. Therefore, we want to distinguish the link densities according to the hierarchies in the network. We do this by decomposing the elements in $P_{AEI}$ further
$i_{X}=\frac{I_{c_{X}}+I_{cp_{X}}+I_{p_{X}}}{\frac{1}{2}n_{X}(n_{X}-1)}\,,$ (2)
where $I_{c_{X}}$ represents the interaction between group $X$’s core nodes, while $I_{cp_{X}}$ denotes the interaction between its core and periphery, $I_{p_{X}}$ denotes the interaction among its periphery nodes, and $n_{X}$ denotes the number of nodes in group $X$. We can now subsitute the corresponding terms of Eq. (2) into Eq. (1), which gives us
$P_{AEI}=(\underbrace{\frac{i_{c_{A}}}{\alpha}}_{\text{elite cohesion}}+\underbrace{\frac{i_{cp_{A}}}{\alpha}}_{\text{mass amplification}}+\underbrace{\frac{i_{p_{A}}}{\alpha}}_{\text{mass cohesion}})+(\frac{i_{c_{B}}}{\alpha}+\frac{i_{cp_{B}}}{\alpha}+\frac{i_{p_{B}}}{\alpha})-\underbrace{\frac{2\times e_{AB}}{\alpha}}_{\text{bridge}}\,,$ (3)
where $\alpha$ is the normalization factor equal to $(i_{c_{A}}+i_{cp_{A}}+i_{p_{A}})+(i_{c_{B}}+i_{cp_{B}}+i_{p_{B}})+2\times e_{AB}$.
We call this decomposition polarization spectrum as it allows us to further study the direct effects of different groups and hierarchies on the polarization score. By splitting the observed polarization score into components, we can detect the differences between groups and their hierarchies on the polarization measure and determine whether their contributions have changed over time. When we say ’contribution’ or ’impact’ of a certain group on observed polarization, we specifically refer to these components. This is valuable information in the context of polarization, where we are often interested in monitoring the evolution of the opposing groups and their power dynamics.
Introducing hierarchical groups also enables us to operationalize the polarization processes. We call $i_{c_{X}}/\alpha$ the elite cohesion of group $X$. The higher the value, the more organized is the core. Similarly for $i_{p_{X}}/\alpha$, which we call mass cohesion of group $X$. Structurally, this value quantifies the centralization of leadership in a given group. Lower values indicate more centralized leadership, as most links would be directed towards amplifying the core rather than connecting with peers in the periphery. Finally, we name $i_{cp_{X}}/\alpha$ as mass amplification, capturing the intuition that surrounding periphery act as an amplifying body for its group’s core voices.
We expect the hierarchy of users playing a significant role in terms of polarization. To better understand the impact on score, we also measure the marginal polarization for both core and periphery nodes. We define marginal polarization as the change in structural polarization when a node of a specific type is added to the network, which enables us to assess how introducing an average elite user versus an average non-elite user affects the network’s overall polarization.
A core node in group $A$ has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ external links to group $B$. For a periphery node belonging to group $A$, it can have $k_{p_{A}}$ links to other periphery nodes. It amplifies $k_{cp_{A}}$ core nodes and connects to $k_{out}$ links outside its own group. By substituting the average of observed values into these “decomposed” degrees, we can approximate the marginal polarization for a mean core node as
$\Delta_{core_{A}}P_{AEI}\approx 2\times\frac{\frac{\langle k_{c_{A}}\rangle+\langle k_{cp_{A}}\rangle}{(n_{A})^{2}}-\frac{\langle k_{out_{B}}\rangle}{n_{A}n_{B}}}{\alpha}\,,$ (4)
where $n_{X}$ denotes the number of nodes in group $X$. The formula is equivalent for the situation where an average periphery node is added to the polarized system $(\Delta_{periphery_{A}}P_{AEI}$), but you need to substitute the $k$ values with those reflecting the link distribution of peripheral nodes. Interestingly, the formula suggests that having an equal number of links in and out is not enough for having zero net impact on polarization when group $A$ is smaller than group $B$. See A.6 for full derivation of this marginal polarization formula.
In addition to structural polarization, we aim to quantify the degree of issue alignment separately for the elite and mass. Issue alignment contributes to political polarization by creating a situation where individuals or groups align themselves strictly along specific issues. When people strongly identify with a particular issue or group and use it as a defining factor for their political stance towards other topics, it can lead to several detrimental consequences [CSG+21]. We measure alignment using the normalized mutual information (NMI) method [CSG+21] for cores and peripheries separately. In essence, it quantifies the extent to which knowing the group of an user in one topic informs us about the users’s group in another topic.
Let $G_{1}=(V_{1},M_{1})$ and $G_{2}=(V_{2},M_{2})$ denote networks representing issues 1 and 2, respectively. For estimating each user’s stance on a specfic issue, we use their group membership in the given network. Since this assessment requires user’s presence in both networks, we add the stances for each node in $V_{1}\cap V_{2}$ in stance vectors, $s_{1}$ and $s_{2}$, corresponding to the respective issues, i.e. these stance vectors effectively encode each user’s group membership in the networks. Finally, we compute the issue alignment separately for the distinct hierarchies; NMI($s_{1}^{c}$, $s_{2}^{c}$) for the cores (elite alignment) and NMI($s_{1}^{p}$, $s_{2}^{p}$) for the peripheries (mass alignment). This measure is constrained between 0 and 1, where 0 indicates no alignment, and 1 indicates complete alignment between the issues in question.
3 Hierarchical polarization in Finnish Parliamentary Elections 2019 and 2023
Topic $|N|_{19}$ $|E|_{19}$ $|N|_{23}$ $|E|_{23}$ $AEI_{19}^{*}$ $AEI_{23}^{*}$
Climate politics 16 639 50 605 11 264 34 509 0.45 0.53
Immigration politics 6 739 20 923 9 519 36 889 0.50 0.61
Social policies 13 926 39 778 11 395 34 173 0.36 0.49
Economic policy 5 880 10 962 10 031 37 111 0.28 0.57
Education policy 9 843 20 314 12 814 35 619 0.29 0.47
Table 1: Retweet networks were constructed for both election years, encompassing five distinct networks each. These networks were constructed based on Twitter data obtained over a span of 12 weeks leading up to the respective election day, which were 14.4. for 2019 and 2.4. for 2023. $|N|$ represents the count of unique nodes (users), and $|E|$ denotes the count of unique edges (retweets) in network after the preprocessing. $AEI^{*}$ denotes the normalized structural polarization computed according to [SCK22], where AEI stands for Adaptive EI-index. Last column of the table shows how polarization has increased significantly for all topics studied here.
3.1 Data
We use data collected from Twitter (Public API v1 and v2) during the Finnish parliamentary elections in 2019 and 2023. To ensure consistency when comparing these two snapshots, we consider a 12-week period leading up to the election day, resulting in the following periods: from January 21, 2019, to April 14, 2019 for the first elections, and from January 9, 2023, to April 2, 2023 for the second elections. For each year, five networks are constructed from sets of keywords related to larger topics, such as immigration and climate change (see Appendix in [CSG+21] for more details). Before constructing the networks, retweets that contain keywords belonging to two or more topics are removed from the dataset, as they would produce a systematic error, i.e. inflate the extent of alignment between networks.
The largest connected component is extracted for each network, as done in [GMGM18, CSG+21, SCK22], with the subsequent removal of self-loops and parallel edges, as performed in [SCK22, CSG+21]. Finally, a network can be viewed as a graph where each user contributing to the topic is assigned to one node. In this graph, an edge between two nodes represents a retweet, which can be deemed as an agreement or a shared point of view on a selected issue between the corresponding users [GMGM18, CSG+21, SCK22]. Basic information about each network can be found in Table 1, and a summary of the network partition outputs in Table 2.
We tested the assumptions stated in the previous section independently: To verify the presence of political figures within cores, we assessed the probability of a politician being found in the core compared to the periphery. We found that politicians who ran for elections are at least 3-4 times more likely to be in the core than in the periphery (see A.1 for more details). To view the connections between the outer periphery and the core nodes as mass amplification, we confirmed this assumption by checking that the majority of retweets flow towards the core nodes (see A.2 for more details). Thus, we find our assumptions reasonable.
Table 1 demonstrates the drastic increase in structural polarization for all networks from 2019 to 2023. Over these four years, normalized AEI more than doubled for the economy network. The climate network showed the smallest percentage increase in polarization, approximately 18 %. For the immigration network, AEI increased from 0.50 to 0.61, reaching the highest polarization value observed among our networks. This extent of development in polarization over such a short period made us investigate its structure in more detail.
Assortative Hierarchy
Topic $G_{19}$ $G_{23}$ $A_{19}$ $B_{19}$ $A_{23}$ $B_{23}$
Climate politics
Immigration politics
Social policies
Economic policy
Education policy
Table 2: Each network studied here has an assortative structure and each identified polarized group has some type of hierarchy that could not be explained with Erdős–Rényi model. We confirm assortativity if the two-group model yields a lower description length than the single-group model. Similarly, hierarchy within each group is confirmed if the CP-model results in a lower description length compared to the null model.
3.2 Groups have unequal impact on the overall structural polarization
Figure 3: Polarization spectrum for structural contributions of different groups and their hierarchies to AEI-score. The figure illustrates the predominant influence of elite cohesion ($i_{{c}_{A}}$ $i_{{c}_{B}}$), mass amplification ($i_{{cp}_{A}}$ $i_{{cp}_{B}}$), and mass cohesion ($i_{{p}_{A}}$ $i_{{p}_{B}}$) on the overall score. The green part of spectrum represents the impact of the bridge between the opposing entities ($2\times e_{AB}$). The contributions of different hierarchical members not only vary within individual networks but also across the distinct networks. The part of the spectrum that corresponds to the internal structures is shifted to the left by an amount equal to the cross-interactions. This enables us to read the unnormalized AEI score for each network directly from the figure.
We encapsulated the contributions of different components into a single plot (see Figure 3), where separate spectrums are composed for each year separately for all topics. Here, groups labeled A (in red) refer to left-leaning groups, while groups labeled B (in blue) refer to right-leaning groups, respectively (see A.3 for labeling details). This allows us to track the changes in the dynamics of left-right groups and their embedded hierarchies in the polarized milieu online.
Results show many interesting findings. First, the influence of polarized groups on the observed structural polarization varies widely. In fact, in no network was the influence of both groups close to balanced, even when the groups were of similar sizes. In 2019, left-leaning groups contributed the most to the structural polarization in the climate, social security, and economic policy networks, whereas right-leaning groups were more dominant in the immigration and education networks. However, in 2023, right-leaning groups accounted for the majority of the observed polarization in 4 out of 5 networks. The left-leaning group remained the more contributing group in the economic policy network, although its share became slightly smaller.
The polarization spectrums also reveal the influence and evolution of distinct hierarchical groups. For instance, in the immigration network, the most polarized topic in both years, the elite cohesion of the right-leaning group had a relatively large impact on polarization. The same elite were also significantly more organized than those in the left-leaning group, although the elite cohesion of the latter increased in 2023. The opposite phenomenon was observed for economic policy network, where the left-leaning group’s elite cohesion was higher, while right-leaning group elite evolved into more organized structure in 2023. Note, how in all networks, at least one of the elite groups had become more cohesive during the four years; Right-leaning elite became more cohesive in four networks and left-leaning elite in three networks, indicating a stronger role of elite in certain topics. The most drastic swap in terms of contributions of a hierarchical groups was witnessed in the mass amplification patterns in climate network, where right-leaning public outrun the previously dominating left-leaning public. Although in the absolute terms, the publics’ interaction patterns tend to explain most of the observed polarization in all networks, the relative contribution (when compared to the sizes of these hierarchical groups) of the elite members is, on average, greater, which we will see later in Figures 4 & 5B.
Communication between polarized groups has mostly dropped or remained at the same level across all topics. This type of communication, labeled as bridge, is the only interaction that could structurally decrease polarization. Its share of the spectrum has shrunk in all networks except the immigration network.
To summarize the changes in power dynamics between the two groups, we visualize the polarization-increasing components of AEI alongside the size distributions of groups in a more compact format (see Figure 4). For each network and year, we plot a thicker rectangle to illustrate the fraction each group (including their hierarchical groups) contributed to the observed AEI. A thinner bar shows the relative sizes of each group. The lower rectangles correspond to the data from 2019, while the upper rectangles represent the results from 2023. These rectangles, which depict the distribution of contributions, sum up to one. In a balanced scenario, the red and blue shaded areas would be symmetrical along the vertical line.
In the context of immigration, the overall impact of the right-leaning group exceeded that of the left-leaning group in both years, despite the right-leaning group being smaller in size. A drastic change was observed in the climate network, where in 2019, the left-leaning group was more dominant. However, after four years, the right-leaning group managed to organize its elite and public components more effectively, resulting in a greater impact on polarization, despite left-leaning group being larger in 2023. The plot for the economy network also illustrates how the size of a group does not necessarily correlate with its impact on the polarization score. In both years, the right-leaning groups were significantly larger, yet the left-leaning groups accounted for a larger portion of the observed polarization.
Three main findings from the evolution graphs (Fig. 4) stand out: first, the increase in structural polarization does not follow uniform patterns, second, a larger group size does not always lead to a higher contribution to the observed polarization, and third, the elite play relatively higher role than the mass. The networks shown are the most polarized ones, yet the underlying contributions of the groups are very different. We see, for instance, how right-leaning group compensated its smaller size with more organized core and greater mass amplification in both years, leading to significantly higher impact on polarization when compared to the impact of left-leaning group. The greater relative contribution of the elite members is generally evident across all networks when reflecting it to the size of their groups shown in the thinner bars. We also confirm this by computing the marginal polarization separately for the elite and mass individuals in Figure 5.
Figure 4: Evolution of power structures in most polarized networks: Contributions of different structures to overall polarization is shown in the thicker bars for the most polarized networks. Mass amplification and cohesion are combined here into one color. Thinner bars illustrate the size distributions of both elite and mass groups within the networks. The upper set of bars corresponds to the results for 2023, while the lower set represents those for 2019. Each bar sums up to 1. These illustrations demonstrate visually how the increases in structural polarization does not follow uniform patterns, larger group sizes do not always lead to higher contributions to observed polarization, and the elite play a relatively higher role than the masses. The similar plots for the remaining topics can be found in Appendix A.5.
3.3 Elite is significantly more aligned than the mass
Figure 5: A: Elites are consistenly more aligned than masses across all topic pairs. Elites became more aligned in 2023, together with a smaller increase in the alignment of mass opinion on various issues. To capture the uncertainty around the observed values, we bootstrapped 500 pairs of networks for each topic pair. Each bootstrap sample represents a perturbed version of the original network, where the sizes of cores and peripheries are subject to random fluctuations. B: Elite tend to have higher marginal polarization as well compared to the mass. In both years, adding a new elite member to the economy network had the greatest impact. A weighted average of both groups’ marginal values is applied to obtain a single value representing the hierarchy’s mean effect on AEI.
Elite members are consistently more aligned than the public in all topics studied here, as shown in Figure 5A. Higher alignment suggests that the elite regularly maintain more predictable stances on various political themes, as typically expected. In 2019, among both elite and mass, topic pairs immigration & education and immigration & climate were the most aligned. Conversely, the lowest alignment was found between the topics immigration & social and education & social.
After four years the gap between elite and mass alignment widened – not only did the elite show a much higher level of issue alignment (the mean level doubled), but the disparity between the mass and the elite also increased in 2023. The overall level of mass alignment did not stay at the same level as in 2019 either. Instead, the mass also experienced a notable increase in alignment across all topics, as the mean line, corresponding to the average of alignment over all the topics, rose from 0.13 to 0.44. For the elite groups, eight out of ten pairs of topics had an NMI value over 0.8, which can be considered very high. Results show that knowing elite member’s stance on climate politics would reveal the most information on their views about immigration policies compared to the other topics included in this study. Interestingly, climate & economy experienced the smallest increase in alignment.
To evaluate deeper the relationship between the elite and mass alignment, we employ bayesian linear regression. As our study contains five distinct topics, the total number of pairs are ten for each year, resulting in a total of twenty samples for the entire period. We fit a linear regression using elite alignment value as an independent variable and mass alignment as dependent variable (see Figure 6). The bayesian framework enables the quantification of uncertainty in model parameters, which we find crucial due to the small sample size. While we acknowledge that our approach may oversimplify the relationship between the elite and mass, we believe that our analysis sheds light on the behavior of these distinct hierarchical groups and highlights how the interdependence between the variables has intensified over time. To be more specific, here we consider the slope of the regression to correspond to the strength of the relationship between the elite and mass alignment. It’s also worth reminding that the linear dependence between the variables we are about to report doesn’t necessarily imply causality. However, the slope of linear regression is tightly connected to the pearson correlation coefficient, as
$\text{slope}=r\times\frac{\text{standard deviation of }y}{\text{standard deviation of }x}\,,$ (5)
where slope is the slope estimate of the fitted regression line and $r$ is the pearson correlation coefficient.
The posterior distribution for the slope $\beta$ is centered at 0.607 (95% CI [0.60, 0.61]), with the model’s intercept being -0.08. Hence, the expected slope, and consequently, the correlation, is positive. This suggests that as elite alignment increases, mass alignment also increases. It’s evident from the regression results that the elite tend to be more politically aligned than the general public, which aligns with common expectations. However, intriguingly, the model also offers an estimate for the “baseline alignment”, i.e. the degree of alignment among the elite even when the mass shows no alignment whatsoever. According to our model, the point estimate for the baseline alignment is approximately 0.13.
Figure 6: Each data point represents the observed alignment between pair of topics separately for the elite (x-axis) and public (y-axis). To assess and compare the strength of the relationship between these variables, bayesian linear models were fitted. The subplot illustrates the posterior distribution of the slope parameter. The linear correlation is evidently positive between the variables, and for most topics in 2023, the mass exhibits higher issue alignment than the expected posterior mean.
4 Conclusion
4.1 Discussion
This study investigated whether polarized groups and their respective hierarchies have equal influence on polarization observed in online systems. We proposed a method to identify these structures within a single social system and analyzed their different impacts on two aspects of polarization: structural polarization and issue alignment-based polarization. We illustrated the use of the method by applying it to Finnish polarized networks.
Our findings revealed that polarized groups had an unequal impact on overall structural polarization across all the networks studied, indicating lack of balanced systems where both opposing groups would contribute equally to the polarization score. Specifically, we found that left-leaning groups played a more dominant role in three out of five topics (climate, social, and economy) in 2019. However, by 2023, the right-leaning group became the dominant force in both climate and social politics, mostly driven by a more organized elite and greater mass amplification. Finland elected a left-leaning government in 2019 and right-leaning government in 2023 [GS23].
The hierarchy of the user in a polarized networks matters. Not only did we find that the elite consistently aligned more on distinct political topics than the mass within our networks, but we also observed that the elite had a relatively higher impact on structural polarization, as evidenced by their respective marginal values. Additionally, we discovered that all networks, which experienced an overall increase in polarization during the study period, had at least one elite group that increased its cohesion. This reminds us that a network’s increased polarization does not necessarily depend on greater public amplification. Rather, it suffices for the elite to become more organized, as exemplified for the right-leaning group in the education network.
Elite are significantly more aligned than the public online, and the alignment intensified over the four years across all the topics. The absolute difference in the average level of elite alignment more than doubled over the same time span. We asked whether the public online is somehow immune to the increasingly polarized environment. Although this study cannot respond, whether the drastic increase in the mass alignment is due to the elite being more polarized, our results clearly suggests that certain ideological realignment has taken place among the public too. The relative increase in the issue alignment was more pronaunced for the mass (+228 %) than for the elite (+121 %). We also observed that in 2023, a significant number of observations clustered above the mean for various topics, suggesting a stronger association between elite and mass alignment. Hence, there is evidence that the mass has at least not remained isolated from the trend towards increased alignment.
Again, while our analysis does not definitively pinpoint whether the elite drove this increase in mass alignment, it’s clear that socially influential figures are becoming more politically aligned. This trend could potentially influence the public, aligning with past research highlighting how they often take cues from elite figures. Furthermore, our model suggested that theoretically, the alignment among the elite would persist even if the mass showed no indication of aligning on issues. This could suggest that maintaining predictable stances on a range of political topics serves as a distinguishing feature (or strategy) that sets the elite apart from the general public.
Our contribution does not address the concept of an “ideal polarized network”. Would such a network be non-hierarchical, or would it exhibit balanced contributions from both opposing parties? It is important to acknowledge that some degree of disagreement is not only inevitable but also essential in a healthy democracy. Ideally, disagreement fosters debate, encourages diverse perspectives, and promotes critical thinking, which can be deemed as fundamental components of democratic processes. How these potentially desirable dynamics would be reflected in the network structure is an open question. However, we believe it is crucial to recognize the prevalent imbalances we have in today’s polarized networks and to take them into account in the polarization-related literature. Our study also reveals a concerning trend: a rapid decline in communication patterns and a sharp rise in issue alignment, both which have happened within a relatively short timeframe.
Social media is sometimes seen as a decentralized platform for content creation [TLT+18, LPBF19] that is not monopolized by a few influential entities. However, our observations could challenge this perception. The elite groups consisted of very small number of individuals, and these were literally in the core of the polarized networks, making them structurally dominant. Yet they had greater marginal impact on polarization and were significantly more politically aligned. This raises questions about the true decentralization of social media platforms and whether they genuinely represent the diversity of political voices in our societies.
4.2 Limitations
While our study provides valuable insights into some structural mechanisms of polarization, it is not without limitations.
Many influential figures use Twitter, and it is popular social media platform among the public too. However, while the core-periphery concept has been extensively employed to model hierarchies in networks, and our manual check indicated that a majority of political candidates fell within the identified cores, it’s essential to recognize that these findings may not necessarily extend to the behavior of the “actual” political elite (which is also relatively vague concept) in society. The same applies to the mass, which represented the non-elite nodes in the network. However, it’s worth noting that, as highlighted in previous studies, Twitter users are often perceived as more politically engaged than an average citizen. This limitation is interesting, as it can imply that the mass (and the elite) outside Twitter might be even less aligned than the elite on social media, or the other way around. The function of the public within a polarized network was simplified to act as amplifiers of the messages produced by the cores. While this simplification is not uncommon in similar studies, we believe that the temporal interactions and dynamics between a network’s cores and peripheries are more complex.
While each method presented here could be easily extended to characterize polarization in the case of multiple groups, we have chosen to concentrate on the typical settings involving two groups to facilitate comparisons between the networks.
Considering network ties as directed could enhance the accuracy of network decomposition, as the direction of endorsements is not random. Assigning greater importance to nodes with an in-degree larger than their out-degree may be in some cases desirable. Likewise, when an elite endorses someone who does not reciprocate, it might suggest a special user, such as an organizational account. In this work, we interpretted the edge between two nodes more as a mutual agreement than directed endorsement.
4.3 Future work
We mainly focused on studying structrual differences between the polarized groups and embedded hierarchies in polarized networks. Building upon the insights gained from this study, future research could investigate deeper the content and attitudes within cores, and how they reflect to the structural properties of the network. Political discussions are often perceived as a clash between the status quo and change. Do these respective groups exhibit similar patterns across all networks? Can the structural properties alone of a polarized group offer insights into the positions held by its members? How are these connected to the forces driving polarization online?
Social media platforms have experienced a surge in polarization in recent years, with conflicting findings on whether it is the public or the elite driving this phenomenon. We do not claim that the methods presented here provide a direct answer to this question. However, by conceptualizing the components linked to structural polarization, this study facilitates the examination of temporal changes. Particularly when there are data available from a longer period, it becomes possible to detect persistent structures and identify strengthening groups, whether in terms of mass amplification or elite cohesion. These factors could contribute to a more nuanced understanding of the possible under- or overrepresentation of specific groups. In fact, the introduced polarization spectrums enable us to compare how different social media platforms exhibit disparities in power structures and dynamics among competing political groups. Furthermore, treating elite and mass polarization as distinct processes could assist researchers in detecting suspicious activities that contribute to polarization in online networks—such as abrupt spikes in mass amplification that deviate from their expected trajectory.
The significance of hierarchies and their impacts on online systems should not be underestimated. Our findings enable the deeper exploration into these matters, enriching the comprehension of mechanisms that drive polarization.
Appendix A Appendix
A.1 Political actors in cores versus peripheries
To verify that the presence of political elite is higher in the cores than in peripheries, we computed the following for each network: How much more likely it was for a randomly selected node from the cores to belong to the list of political candidates compared to a node sampled from the peripheries, i.e.
$\gamma=\frac{\text{Probability of core node being a candidate}}{\text{Probability of periphery node being a candidate}}$
Topic $\gamma$
CLIMATE 3.3
IMMIGRATION 2.8
SOCIAL SECURITY 3.1
ECONOMIC POLICY 4.0
EDUCATION 3.6
Table 3: Values of $\gamma$ demonstrate the higher chance of finding a politician from the cores versus peripheries.
A.2 Public Amplification
Figure 7: To view the core-periphery interactions as public amplification, we confirm that the majority of connections are directed towards the core. In all networks examined, most of the links between the core and periphery consist of retweets originating from the outer periphery. This proportion remains relatively consistent across different polarized groups. Therefore, it’s reasonable to characterize this dynamic as public amplification.
A.3 Political parties presence in inferred polarized groups
Figure 8: Majority of the candidates from left-leaning parties are grouped together in the inferred polarized groups, as are those from right-leaning parties.
A.4 Characterics of polarized groups
Group A Group B
Topic C P E Coh. Ampl. P Coh. C P E Coh. Ampl. P Coh. B
CLI ’19 2.2 42.4 16.6 40.1 8.6 2.8 53.0 4.0 16.8 5.2 8.7
CLI ’23 3.6 65.8 6.3 12.9 4.7 1.7 29.0 11.9 52.0 9.1 3.1
IMM ’19 3.7 71.8 2.2 7.6 2.2 2.5 22.1 29.9 50.9 5.2 1.9
IMM ’23 3.7 51.8 9.9 17.6 4.0 3.6 41.0 20.5 39.1 5.7 3.1
SOC ’19 1.5 39.0 9.9 37.1 9.8 4.4 55.1 7.2 18.2 6.7 11.3
SOC ’23 4.2 57.3 8.4 15.6 6.7 2.7 35.8 13.7 42.0 8.7 4.9
ECO ’19 1.1 16.2 8.2 54.2 17.8 4.7 78.0 2.2 10.2 4.3 3.1
ECO ’23 2.4 28.9 23.8 36.0 8.5 4.1 64.7 7.3 17.2 4.1 3.2
EDU ’19 4.9 81.0 2.9 9.1 3.9 1.2 13.0 10.9 49.1 19.8 4.4
EDU ’23 4.0 60.4 5.6 14.2 6.0 2.4 33.2 16.8 45.1 8.3 4.0
Table 4: All values in the table represent percentages. ’C’ indicates the size of the corresponding core, while ’P’ indicates the size of the corresponding periphery, divided by the total number of nodes in the network. For the polarization spectrum components, ’E/P Coh.’ (Cohesion) column reflects either elite or public cohesion, while the ’Ampl.’ (Amplification) column shows the public amplification values. Bridge denotes the component representing the overall inter-communication between groups A and B. See more details in the caption for Figure 3.
A.5 Power dynamics
Figure 9: Evolution of power structures for social and education networks. See caption of Figure 4 for more details.
A.6 Marginal polarization
How much does the overall interactional polarization increase if you added an “average” core (or periphery) node to group A? Assume that an average core node has $k_{c_{A}}$ links to other core nodes in the same group, it receives $k_{cp_{A}}$ links from the periphery and forms $k_{out}$ links to the other group B. The marginal polarization could be written then as
$\Delta P_{AEI}(\text{+ corenode to A})$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}+1}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}}{\alpha^{\prime}}$
$\quad-\frac{\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\binom{n_{A}}{2}}+\frac{(I_{c_{B}}+I_{cp_{B}}+I_{p_{B}})}{\binom{n_{B}}{2}}-2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (6)
where $I_{\bullet}$ is the observed number of links within the corresponding structure, and $E$ denotes the observed number of links between groups $A$ and $B$. Each term is adjusted for density, as described in the main text. We approximate $\alpha^{\prime}=\alpha$. Hence we get for average core node the following marginal:
$\Delta P_{AEI}(\text{+ corenode to A})$ $=\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}+1))}-2\times\frac{E+k_{out_{B}}}{(n_{A}+1)n_{B}}-\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}-1))}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (7)
$\approx\frac{\frac{(I_{c_{A}}+k_{c_{A}}+I_{cp_{A}}+k_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}))}-2\times\frac{E+k_{out_{B}}}{n_{A}n_{B}}-\frac{(I_{c_{A}}+I_{cp_{A}}+I_{p_{A}})}{\frac{1}{2}n_{A}(n_{A}))}+2\times\frac{E}{n_{A}n_{B}}}{\alpha}$ (8)
$=\frac{\frac{(k_{c_{A}}+k_{cp_{A}})}{\frac{1}{2}n_{A}(n_{A}))}-2\times\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$ (9)
$=2\times\frac{\frac{(k_{c_{A}}+k_{cp_{A}})}{(n_{A})^{2}}-\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$ (10)
And for the others:
$\Delta P_{AEI}(\text{+ core node to B})$ $\approx 2\times\frac{\frac{(k_{c_{B}}+k_{cp_{B}})}{(n_{B})^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}}}{\alpha}$ (11)
$\Delta P_{AEI}(\text{+ periphery node to A})$ $\approx 2\times\frac{\frac{(k_{p_{A}}+k_{cp_{A}})}{(n_{A})^{2}}-\frac{k_{out_{B}}}{n_{A}n_{B}}}{\alpha}$ (12)
$\Delta P_{AEI}(\text{+ periphery node to B})$ $\approx 2\times\frac{\frac{(k_{p_{B}}+k_{cp_{B}})}{(n_{B})^{2}}-\frac{k_{out_{A}}}{n_{B}n_{A}}}{\alpha}$ (13)
A.7 Statistical testing
Our analysis requires the network to be decomposed into $k$ polarized groups (assortativity assumption) after which each polarized group is decomposed further into separate hierarchical groups (hierarchy assumption). Therefore, in case of $k=2$, the overall number of different groups would be four. From now, we assume $k=2$, and label the identified polarized groups as $A$ and $B$. For the hierarchical decomposition, we label the cores of $A$ as $c_{A}$ and its periphery as $p_{A}$ (similary for B).
We know that most existing methods for determining the modular structure of networks face issues. They often overlook the statistical evidence supporting the identified patterns, making it difficult to distinguish genuine structures from noise. In our case, we want to be sure about two things: 1) Is there enough evidence that the network can be decomposed into two groups? and 2) If so, is there enough evidence that these groups can be partitioned further into separate hierarchies?
Each network undergoes the following partitioning pipeline: 1) Network is partitioned into two blocks with the non-uniform planted partition model. 2) To find the embedded hierarchies, we consider the groups independently, and run the core-periphery algorithm provided here. We apply the traditional hub-and-spoke description of core-periphery structure as it has been shown to explain online amplification better than the layered one. We run this two-step decomposition for 100 times and save each obtained partition for the statistical significance testing.
We want to be explicit about how we measure significance and what we do with that information. First, if it turns out that the network has no significant assortative groups, then we conclude that there is no evidence for the network being at least structurally polarized. In our experiment, all networks were better described with two groups in terms of description length. We then shift our attention to the inferred hierarchies within each group.
Fist, defining a hierarchy is harder than defining the absence of it. Perhaps the most common baseline model in network science, Erdős–Rényi model, generates random graphs, where each edge $(i,j)$ is included in the graph with probability $p$ independently from every other edge. Hence, the possible heterogeneity that emerge in such graph is purely due to random fluctuations, making it decent candidate for null model.
A stricter null model would be the configuration model, another common option for testing the statistical evidence of an observed structure. Essentially, it is a mathematical model used to generate random networks with a specified degree sequence. Interestingly, previous research has shown how the degree sequence of a network can fully explain the core-periphery structure in certain situations [KM18]. We selected ER-model over the configuration model for three reasons:
  1. 1.
    We use the core-periphery structure to identify the structurally dominant users, and it is acceptable for this significance to stem from a user’s higher degree.
  2. 2.
    Some work (e.g. [YS23]) has discussed the possible shortcomings of using configuration model in detecting core-periphery structures, claiming it likely leading to low power for the statistical test. They illustrate this with an ideal core-periphery structure, for which the same original network is the only one that maintains the fixed degree sequence. Therefore, no rewiring is possible, resulting in a hypothesis test yielding a p-value of 1. [KM18] claims that an additional block, such as a community or another core-periphery, is required to conclude the significance of the observed core-periphery. However, as far as we know, there isn’t a generative model capable of inferring multiple core-periphery structures simultaneously, and such model is typically required to measure the description length.
  3. 3.
    Erdős–Rényi model is an intuitive and simple model to understand, and it does not produce self-loops or parallel edges.
Ideally, we would apply directly a constrained version of nested stochastic block model [Pei14a] because of the hierarchical nature of our overall decomposition approach. Unfortunately, it is out of scope of this work to develop such method, and hence we employ the method described in the main text for model selection.